Find the surface area
1.h=5cm,w=7cm,l=5cm

Answer by engsaoudah(9)

total surface area of a cuboid=2[(length*width)+(width*height)+length*height)]
h=5cm w=7cm l=5cm
2[(5*7)+(7*5)+(5*5]=2[(35+35+25)]=2*95=190cm2